# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

### v3

@@ III-C Loss Functions
Fig. 5: Illustrates the visual effects of operations performed on watermarked images. The original image $I_{o}$ and the watermarked image $I_{w}$ are shown in the first and second rows, respectively. The results of applying different operations to the watermarked image, denoted as $I_{n}$, are displayed in the third row. The three columns on the right showcase the effects of deepfake operations, while the remaining columns present the results of benign operations. The fourth row shows the watermarked residual signals, represented as $|I_{o}-I_{w}|$, which are processed to enhance the visualization of differences. All images are sized $128\times 128$.
@@ IV-A Implementation Details
Fig. 6: More malicious distortions. From top to bottom are the origin image $I_{o}$, the watermarked image $I_{w}$, the noised image $I_{n}$, the residual signal $|I_{o}-I_{w}|$, and $|I_{n}-I_{w}|$. The columns indicate different types of transformations, including SimSwap, GANimation, and various StarGAN attribute modifications (Male, BlondHair, BlackHair, BrownHair, Young). All images are resized to $256\times 256$ for consistency.
Manipulation Pools. In order to comprehensively evaluate the robustness and adaptability of the algorithm under different distortion conditions, this study constructs two types of operation pools: the Common Pool and Malicious Pool. The Common Pool includes a variety of common image distortion operations, such as {Identity, JpegTest, Resize, GaussianBlur, MedianBlur, Dropout, SaltPepper, GaussianNoise} . These operations are used in the training and testing stages to simulate the general distortions that may occur in real-world scenarios, while ensuring that the facial content of the image remains unchanged. The Malicious Pool includes two common types of deepfake technologies: face swap and face replay. Specifically, SimSwap[2] is used for face swap, while GANimation[4] and StarGAN[3] are used for face replay. In the SimSwap task, the target face is randomly selected from the CelebA verification set. ”In the tasks of GANimation and StarGAN, the target expressions and attributes are randomly selected. The attributes are chosen from the set Male, Young, BlackHair, BlondHair, BrownHair. All forged images are generated using the publicly released pre-trained models to ensure the high credibility and diversity of deepfake images.
TABLE I: objective visual quality evaluation of watermarked images on celeba-hq: model, image size, message length, psnr, and ssim
@@
Ours $128\times 128$ 30 47.5555 0.9957
@@
SepMark[18] $256\times 256$ 128 38.5646 0.9328
Ours $256\times 256$ 128 43.3082 0.9893
Parameters. This research is based on PyTorch[46] and runs on an NVIDIA RTX 4090 GPU. The Adam optimizer[47]is initially set with a learning rate of 0.0005, which is later dynamically adjusted during training. The total loss function is composed of multiple loss terms, which are combined by weighting to balance the influence of each loss on model optimization. As shown in Formula (7), during the training process, the weight parameters $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to 0.01, 1, 10, 10, and 10, respectively. These values are based on the preliminary experimental results and the empirical setting of the contribution of each task target to the overall optimization, aiming at achieving a good balance and stable convergence between embedding quality, detection ability and structural consistency.
@@
TABLE II: Robustness evaluation of watermark recovery on CelebA-HQ under common and deepfake distortions at 128$\times$128 and 256$\times$256 resolutions, measured by bit error rate (BER).
$128\times 128$ $256\times 256$
Distortion MBRS[27] CIN[28] PIMoG[37] SepMark[18] Ours FaceSigns[17] SepMark[18] Ours
Tracer Detector Tracer Detector Tracer Detector Tracer Detector
Identity 0.0000% 0.0000% 0.0366% 0.0000% 0.0000% 0.0000% 0.0000% 0.0136% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.2597% 2.7514% 19.5562% 0.2136% 0.2172% 0.0000% 0.0000% 0.8258% 0.0075% 0.0069% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0767% 0.0059% 0.0212% 0.0000% 0.0000% 1.0726% 0.0000% 0.0000% 0.0000% 0.0000%
GaussianBlur 0.0000% 22.7786% 0.1169% 0.0024% 0.0035% 0.0018% 0.0021% 0.1671% 0.0000% 0.0274% 0.0000% 0.0056%
MedianBlur 0.0000% 0.0307% 0.0992% 0.0012% 0.0012% 0.0013% 0.0017% 0.0977% 0.0000% 0.0000% 0.0000% 0.0000%
Dropout 0.0000% 0.0000% 0.4828% 0.0000% 0.0000% 0.0000% 0.0000% 17.5615% 0.0058% 0.0000% 0.0018% 0.0027%
SaltPepper 0.0000% 0.0378% 2.3667% 0.0413% 0.0106% 0.0132% 0.0156% 12.3238% 0.0008% 0.0003% 0.0077% 0.0045%
GaussianNoise 0.0000% 0.0000% 12.7396% 0.7460% 0.8735% 0.0211% 0.0327% 7.0697% 0.0578% 0.0622% 0.0039% 0.0045%
Average 0.0325% 3.1998% 4.4408% 0.1329% 0.1409% 0.0047% 0.0065% 12.8914% 0.0249% 0.0121% 0.0017% 0.0023%
SimSwap 19.3744% 48.5068% 8.6745% 13.8255% 50.8829% 0.0287% 46.7845% 49.9463% 7.9068% 45.9117% 0.0086% 44.5411%
GANimation 0.0000% 0.0000% 0.4802% 0.0000% 36.7938% 0.0000% 46.7879% 45.4172% 0.0020% 43.8524% 0.0000% 44.5431%
StarGAN (Male) 18.3133% 67.2568% 9.2044% 0.1157% 52.6003% 0.0051% 46.7828% 50.2617% 0.0033% 45.4624% 0.0011% 44.5395%
StarGAN (Young) 17.0562% 69.0805% 8.7465% 0.1074% 52.3678% 0.0019% 46.7828% 50.3649% 0.0030% 45.5319% 0.0008% 44.5458%
StarGAN (BlackHair) 19.2233% 58.7913% 10.5312% 0.1416% 49.2434% 0.0022% 46.7845% 50.2576% 0.0019% 45.1299% 0.0022% 44.5403%
StarGAN (BlondHair) 18.4478% 72.9733% 10.3506% 0.1712% 48.6084% 0.0013% 46.7811% 50.9829% 0.0022% 44.6953% 0.0009% 44.5415%
StarGAN (BrownHair) 17.6381% 79.5857% 9.0675% 0.0980% 52.3820% 0.0064% 46.7828% 50.7884% 0.0017% 45.9950% 0.0014% 44.5414%
Average 15.7219% 56.5992% 8.1507% 2.0656% 48.9827% 0.0065% 46.7838% 49.7170% 1.1316% 45.2255% 0.0021% 44.5418%
@@
Baselines. To verify the performance of the proposed method, we selected several representative watermarking and proactive watermarks frameworks as baselines. Although this research belongs to the field of proactive watermarks, due to the limited existing methods and the scarcity of open source codes, we also refer to the robust watermarking method to evaluate its stability and extraction performance. In the watermarking domain, three robust watermarking methods were selected: MBRS [27], CIN [28], and PIMOG [37] (suitable for 128×128 images). In the proactive watermarks domain, we include FaceSigns[16] (for 256×256 images) and SepMark[18] (which supports both 128×128 and 256×256 resolutions). All the other baseline methods use publicly available pre-trained model weights to ensure fairness.
@@ IV-B Visual Quality
To evaluate the visual quality of the image after watermark embedding, this paper adopts two indicators: average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM)[48], which respectively quantify the changes in image signal distortion and structural fidelity, fully reflecting the impact of embedding on the image. TABLEI quantitatively verifies that our method has the best visual quality in all baselines and shows better invisibility. Experimental results show that our method achieves the highest PSNR and SSIM values at resolutions of 128×128 and 256×256. This improvement is primarily attributed to SC-GNN and attention-enhanced embedding. In addition, Fig.5 and 6 show the subjective visual contrast results at different resolutions, respectively, indicating that it is almost difficult to distinguish the watermarked image from the original image under common distortion and deepfake attacks. Combining the objective evaluation results in TABLEI with the subjective visual comparisons in Fig. 5 and 6, it is evident that the proposed method achieves superior imperceptibility in both quantitative metrics and human visual perception.
@@ IV-C Robustness Test
The average bit error rate (BER) is selected as the main evaluation index in the robustness test. In any case, the BER value of Tracer should be close to 0%, and the BER value of the semi-robust Detector should be close to 50% when it encounters malicious distortion. In the case of common distortion, the BER value should be close to 0%.
@@
where $W_{o}$ is the original watermark, $W_{\text{rec}}$ represents the recovered watermark, and $l$ denotes the length of the binary watermark sequence. The BER calculates the ratio of error bits to the total length of the watermark, serving as a direct measure of recovery accuracy under various distortion scenarios.
TABLE III: Robustness test for the image $I_{n}$ under common and malicious distortions on cross-dataset LFW
128 $\times$ 128 256 $\times$ 256
Distortion PSNR = 46.4265 PSNR = 43.7051
Tracer Detector Tracer Detector
Identity 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0027% 0.0013% 0.0031% 0.0011%
Resize 0.0000% 0.0000% 0.0000% 0.0000%
GaussianBlur 0.0046% 0.0038% 0.0052% 0.0064%
MedianBlur 0.0023% 0.0011% 0.0011% 0.0008%
Dropout 0.0065% 0.0096% 0.0023% 0.0032%
SaltPepper 0.0116% 0.0223% 0.0043% 0.0066%
GaussianNoise 0.0319% 0.0428% 0.0087% 0.0072%
SimSwap 0.0352% 47.5363% 0.0057% 46.5418%
GANimation 0.0009% 47.5368% 0.0000% 46.5432%
StarGAN (Male) 0.0047% 47.5313% 0.0007% 46.5461%
StarGAN (Young) 0.0038% 47.5356% 0.0012% 46.5431%
StarGAN (BlackHair) 0.0096% 47.5361% 0.0008% 46.5432%
StarGAN (BlondHair) 0.0029% 47.5368% 0.0014% 46.5434%
StarGAN (BrownHair) 0.0072% 47.5363% 0.0007% 46.5468%
Average Common 0.0075% 0.0101% 0.0031% 0.0032%
$\uparrow$0.0028$\uparrow$0.0036$\uparrow$0.0014$\uparrow$0.0009%
Average Malicious 0.0092% 47.5342% 0.0015% 46.5439%
$\uparrow$0.0027—- $\downarrow$0.0007—-
@@
TABLE IV: ablation results for psnr, ssim, and tracer’s bit error rate (%), where ‘×’ denotes non-usage and ‘$\sqrt{}$’ denotes usage.
@@
PSNR $\uparrow$ 45.2321 43.7024 49.0436 42.8801 47.5555
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9517 0.9957
Identity 0.0000% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0000% 0.0000% 0.0198% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0057% 0.0000% 0.0000%
GaussianBlur 0.0202% 0.0021% 0.0029% 0.0021% 0.0018%
MedianBlur 0.0021% 0.0017% 0.0011% 0.0005% 0.0013%
Dropout 0.0008% 0.0000% 0.0000% 0.0000% 0.0000%
SaltPepper 0.0048% 0.0033% 0.0428% 0.0226% 0.0132%
GaussianNoise 0.0032% 0.0208% 0.5491% 0.0108% 0.0211%
Average 0.0039% 0.0035% 0.0777% 0.0045% 0.0047%
SimSwap 0.2175% 0.2139% 0.9971% 0.0713% 0.0287%
GANimation 0.0015% 0.0000% 0.0038% 0.0013% 0.0000%
StarGAN (Male) 0.0247% 0.0067% 0.1207% 0.0069% 0.0051%
StarGAN (Young) 0.0138% 0.0032% 0.1069% 0.0044% 0.0019%
StarGAN (BlackHair) 0.0103% 0.0017% 0.1128% 0.0031% 0.0022%
StarGAN (BlondHair) 0.0119% 0.0027% 0.1249% 0.0063% 0.0013%
StarGAN (BrownHair) 0.0076% 0.0089% 0.1416% 0.0097% 0.0064%
Average 0.0410% 0.0339% 0.2297% 0.0147% 0.0065%
TABLE V: ablation results for psnr, ssim, and detector’s bit error rate (%), where ‘×’ denotes non-usage and ‘$\sqrt{}$’ denotes usage.
@@
PSNR $\uparrow$ 45.2321 43.7024 49.0436 42.8801 47.5555
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9517 0.9957
Identity 0.0000% 0.0000% 0.0000% 0.0000% 0.0000%
JpegTest 0.0000% 0.0000% 0.0152% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0224% 0.0000% 0.0000%
GaussianBlur 0.0017% 0.0027% 0.0064% 0.0013% 0.0021%
MedianBlur 0.0011% 0.0024% 0.0019% 0.0009% 0.0017%

### v4

@@ III-C Loss Functions
Fig. 5: Illustrates the visual effects of operations performed on watermarked images. The original image $I_{o}$ and the watermarked image $I_{w}$ are shown in the first and second rows, respectively. The results of applying different operations to the watermarked image, denoted as $I_{n}$, are displayed in the third row. The fourth row shows the watermarked residual signals, represented as $|I_{o}-I_{w}|$, which are processed to enhance the visualization of differences. All images are sized $128\times 128$.
@@ IV-A Implementation Details
Fig. 6: More malicious distortions. From top to bottom are the origin image $I_{o}$, the watermarked image $I_{w}$, the noised image $I_{n}$, the residual signal $|I_{o}-I_{w}|$, and $|I_{n}-I_{w}|$. All images are resized to $256\times 256$ for consistency.
Manipulation Pools. In order to comprehensively evaluate the robustness and adaptability of the algorithm under different distortion conditions, this study constructs two types of operation pools: the Common Pool and Malicious Pool. The Common Pool includes a variety of common image distortion operations, such as {JpegTest, Resize, GaussianBlur, MedianBlur, Dropout, SaltPepper, GaussianNoise} . These operations are used in the training and testing stages to simulate the general distortions that may occur in real-world scenarios, while ensuring that the facial content of the image remains unchanged. The Malicious Pool includes two common types of deepfake technologies: face swap and face replay. Specifically, for face swapping, we employ SimSwap[2], Uniface[5], CSCS[4], HifiFace[6], and FSRT[3]. These methods span various architectural paradigms, including identity injection (SimSwap), latent space alignment based on StyleGAN (Uniface), explicit pseudo-input supervision (CSCS), 3D-aware face modeling (HiFiFace), and transformer-based disentanglement of appearance, pose, and expression (FSRT). For face reenactment, we adopt GANimation[8] and StarGAN[7], which manipulate facial expressions and attributes through expression-guided deformation and multi-domain attribute control, respectively. In the face swap task, the target identity is randomly selected from the CelebA verification set. For reenactment models, the target expressions and attributes are randomly sampled from a predefined attribute set: Male, Young, BlackHair, BlondHair, BrownHair. All forged images are generated using the publicly released pre-trained models to ensure the high credibility and diversity of deepfake images.
TABLE I: Evaluation of Watermarked Image Quality on CelebA-HQ Using PSNR and SSIM Across Different Models, Image Resolutions, and Message Lengths
@@
ARWGAN[35] $128\times 128$ 30 38.5746 0.9733
EditGuard [45] $128\times 128$ 30 38.1664 0.9796
LampMark[25] $128\times 128$ 30 44.6251 0.9921
Ours $128\times 128$ 30 47.5121 0.9957
SepMark[22] $256\times 256$ 128 38.5646 0.9328
ARWGAN[35] $256\times 256$ 128 31.3319 0.9339
@@
EditGuard[45] $256\times 256$ 128 37.8254 0.9512
LampMark[25] $256\times 256$ 128 42.7825 0.9869
Ours $256\times 256$ 128 43.2530 0.9909
Parameters. This research is based on PyTorch and runs on an NVIDIA RTX 4090 GPU, with the random seed fixed to 42 to ensure reproducibility. The Adam optimizer[52] is initialized with a learning rate of $5\times 10^{-4}$, which is decayed stepwise to $5\times 10^{-5}$ and $1\times 10^{-6}$ at the 3rd and 7th epochs, respectively, to stabilize training. The model is trained with a batch size of 32 for 10 epochs, as preliminary experiments show that performance converges around the 8th–10th epoch, and longer training yields limited improvements while increasing the risk of overfitting. The embedding strength is fixed at $\alpha=0.04$. The total loss function is composed of multiple loss terms, which are combined by weighting to balance the influence of each loss on model optimization. As shown in Formula (7), the weight parameters for each loss term during training are set to $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ as 0.01, 1, 10, 10, and 10, respectively. To verify the appropriateness of these weights, we conducted a hyperparameter sensitivity analysis, as illustrated in Fig. 7. This figure shows the impact of varying each $\lambda$ on the bit error rate. The black star markers denote the adopted settings, each yielding near-optimal performance for its corresponding loss component. These values are determined based on preliminary experimental results and empirical observations, reflecting the relative contribution of each task objective to the overall optimization. The goal is to achieve a well-balanced trade-off and stable convergence among embedding quality, detection performance, and structural consistency.
@@
TABLE II: Robustness evaluation of watermark recovery on CelebA-HQ under common and deepfake distortions at $\times128$ resolution, measured by bit error rate (BER).
Distortion MBRS[31] CIN[32] PIMoG[42] ARWGAN[35] EditGuard[45] LampMark[25] SepMark[22] Ours Tracer Detector Tracer Detector JpegTest 0.2597% 2.4317% 18.4971% 6.7001% 1.4152% 0.0000% 0.2136% 0.2172% 0.0000% 0.0000% Resize 0.0000% 0.0000% 0.0612% 0.0121% 0.0017% 0.0000% 0.0059% 0.0212% 0.0000% 0.0000% GaussianBlur 0.0000% 19.3915% 0.2312% 4.2970% 1.6027% 0.2104% 0.0024% 0.0035% 0.0021% 0.0032% MedianBlur 0.0000% 0.0325% 0.0992% 3.0027% 1.6212% 0.1023% 0.0012% 0.0012% 0.0011% 0.0019% Dropout 0.0000% 0.0000% 0.4713% 0.6905% 0.9257% 0.0000% 1.9125% 0.0000% 0.0000% 0.0000% SaltPepper 0.0000% 0.0364% 2.1571% 0.0000% 0.0925% 5.1365% 0.0413% 0.0106% 0.0187% 0.0211% GaussianNoise 0.0000% 0.0000% 11.4598% 6.3725% 1.6772% 0.2561% 0.7460% 0.8735% 0.0253% 0.0361% Average 0.0371% 3.1274% 4.7110% 3.0107% 1.0480% 0.8150% 0.4176% 0.1610% 0.0067% 0.0089% FSRT 2.4113% 3.2072% 5.0021% 4.3410% 5.7701% 2.9307% 0.7814% 48.3815% 0.0785% 46.7852% SimSwap 24.6044% 40.0828% 7.6547% 46.6012% 45.3202% 0.1975% 20.0295% 50.8829% 0.1637% 46.7815% CSCS 10.1000% 0.2904% 0.2512% 6.2905% 0.9901% 2.3015% 0.6801% 48.7418% 0.0091% 46.7879% HifiFace 2.7251% 7.4510% 1.0211% 14.2084% 6.7104% 2.9713% 0.9904% 48.2511% 0.0081% 46.7872% Uniface 0.4802% 11.8027% 2.7854% 26.2814% 9.1702% 6.2801% 0.3430% 48.9112% 0.0052% 46.7871% GANimation 0.0000% 0.0000% 0.4501% 0.5201% 0.7152% 1.0254% 0.0000% 36.7938% 0.0000% 46.7873% StarGAN (Male) 15.4416% 56.9345% 9.3444% 36.7825% 7.6204% 17.5317% 0.1125% 52.7725% 0.0065% 46.7869% StarGAN (Young) 17.3462% 61.0875% 8.8415% 36.4812% 8.1745% 17.1141% 0.1560% 52.7907% 0.0071% 46.7875% StarGAN (BlackHair) 19.1083% 58.7913% 9.1392% 36.2510% 7.3674% 18.5142% 0.1346% 48.2134% 0.0032% 46.7876% StarGAN (BlondHair) 18.1408% 54.1533% 11.3514% 36.1451% 9.2541% 17.3256% 0.1172% 48.4184% 0.0045% 46.7876% StarGAN (BrownHair) 17.6082% 69.5127% 8.9910% 35.9919% 8.1254% 17.4425% 0.0997% 48.1250% 0.0064% 46.7872% Average 11.6332% 33.0284% 5.8938% 25.4449% 9.9289% 9.4213% 2.1313% 48.3893% 0.0266% 46.7848%
@@
Baselines. To verify the performance of the proposed method, we selected several representative watermarking and proactive watermarking frameworks as baselines. Although this work focuses on proactive watermarking, due to the limited existing methods and scarce open-source codes, robust watermarking methods are also included to evaluate stability and extraction performance. In the watermarking domain, we select five robust methods: MBRS [31], CIN [32] and PIMOG [42] (all only for 128×128), as well as ARWGAN [35], and EditGuard [45]. For proactive watermarking, we include FaceSigns [20] (only for 256×256), SepMark [22], and LampMark [25]. Due to differences in resolution compatibility, each method is reported only in the tables corresponding to its applicable settings. All the other baseline methods use publicly available pre-trained model weights to ensure fairness.
@@ IV-B Visual Quality
To evaluate the visual quality of the image after watermark embedding, this paper adopts two indicators: average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM), which respectively quantify the changes in image signal distortion and structural fidelity, fully reflecting the impact of embedding on the image. TABLE I quantitatively verifies that our method has the best visual quality in all baselines and shows better invisibility. Experimental results show that our method achieves the highest PSNR and SSIM values at resolutions of 128×128 and 256×256. This improvement is primarily attributed to SC-GNN and attention-enhanced embedding. In addition, Fig.5 and 6 show the subjective visual contrast results at different resolutions, indicating that the watermark image is still highly consistent with the original image after embedding, and it is almost difficult to detect the visual difference. Combining the objective evaluation results in TABLE I with the subjective visual comparisons in Fig. 5 and 6, it is evident that the proposed method achieves superior imperceptibility in both quantitative metrics and human visual perception.
Fig. 7: Hyperparameter sensitivity analysis of loss Weights ($\lambda_{1}$$\lambda_{5}$) on bit error rate.
@@ IV-C Robustness Test
The Bit Error Rate (BER) is selected as the main evaluation index in the robustness test. The BER calculates the ratio of error bits to the total length of the watermark, serving as a direct measure of recovery accuracy under various distortion scenarios.
@@
where $W_{o}$ is the original watermark, $W_{\text{rec}}$ represents the recovered watermark, and $l$ denotes the length of the binary watermark sequence. In any case, the BER value of Tracer should be close to 0%, and the BER value of the semi-robust Detector should be close to 50% when it encounters malicious distortion. In the case of common distortion, the BER value should be close to 0%. This 50% threshold is not a strict rule but a practical choice, as our method naturally exhibits a clear gap between the highly robust Tracer and the semi-robust Detector. Such contrast makes the detection process intuitive and effective without requiring parameter tuning. A similar threshold-based design is also adopted in SepMark[22]. Tables II and III provide a quantitative comparison of BER performance across different methods under both common distortions and deepfake forgeries, evaluated at 128×128 and 256×256 resolutions, respectively. Experimental results show that our method achieves the lowest BER in all test scenarios, showing excellent robustness. Under ordinary distortion, the BER of Tracer and Detector are both close to 0%, which ensures that the watermark can be recovered with high precision under all kinds of interference.
TABLE III: Robustness evaluation of watermark recovery on CelebA-HQ under common and deepfake distortions at $\times256$ resolution, measured by bit error rate (BER).
Distortion ARWGAN[35] FaceSigns[20] EditGuard[45] LampMark[25] SepMark[22] Ours Tracer Detector Tracer Detector JpegTest 6.0034% 0.6298% 3.5801% 0.0000% 0.0071% 0.0071% 0.0000% 0.0000% Resize 0.2501% 1.3654% 0.2815% 1.5809% 0.0000% 0.0000% 0.0000% 0.0000% GaussianBlur 7.6214% 0.1971% 3.3701% 0.3151% 0.0000% 0.0298% 0.0000% 0.0056% MedianBlur 3.1214% 0.0997% 3.8504% 0.1825% 0.0000% 0.0000% 0.0000% 0.0000% Dropout 0.7285% 18.4514% 0.9752% 2.4312% 0.0061% 0.0000% 0.0019% 0.0027% SaltPepper 0.5011% 11.4701% 0.1210% 5.2514% 0.0011% 0.0009% 0.0096% 0.0045% GaussianNoise 6.2471% 8.1637% 1.0274% 0.8254% 0.0525% 0.0682% 0.0045% 0.0045% Average 3.4961% 5.7682% 1.8865% 1.5124% 0.0095% 0.0151% 0.0023% 0.0025% FSRT 4.9701% 49.1201% 37.1724% 3.2541% 9.2103% 45.2127% 0.0091% 44.5118% SimSwap 41.5726% 49.3975% 47.5625% 0.7125% 7.9125% 45.9127% 0.0096% 44.5119% CSCS 14.8215% 49.6731% 1.6173% 5.1402% 2.3514% 45.9117% 0.0081% 44.5153% HifiFace 11.2041% 49.7431% 45.0121% 9.1074% 14.0813% 45.8724% 0.0076% 44.5125% Uniface 21.8204% 49.6547% 49.4123% 7.0297% 18.9325% 45.9117% 0.0086% 44.5157% GANimation 4.3051% 45.3601% 7.5911% 0.0165% 0.0020% 43.8325% 0.0000% 44.5131% StarGAN (Male) 32.5146% 50.5931% 2.1241% 16.1040% 0.0073% 45.4375% 0.0013% 44.5123% StarGAN (Young) 32.1901% 50.4131% 3.5414% 16.2167% 0.0030% 45.3125% 0.0009% 44.5127% StarGAN (BlackHair) 31.9904% 50.2376% 2.9812% 15.4211% 0.0016% 45.3644% 0.0032% 44.5178% StarGAN (BlondHair) 32.0541% 50.9129% 2.8574% 16.0487% 0.0021% 44.6956% 0.0011% 44.5131% StarGAN (BrownHair) 32.5706% 50.7404% 3.0145% 16.4575% 0.0037% 45.6759% 0.0021% 44.5179% Average 23.6376% 49.6223% 18.4442% 9.5917% 4.7734% 46.2854% 0.0047% 44.5066%
However, the BER of the contrast method is obviously higher under various common distortion conditions, which indicates that its robustness is weak. Under the deepfake forgery, the BER of Tracer is still close to 0%, while the BER of semi-robust Detector is maintained close to 50%, which effectively distinguishes the true and forged contents. This advantage mainly stems from our combined design of frequency-domain embedding and structural modeling. Existing methods often perform poorly when facing semantic-level transformations, making watermarks more vulnerable. For example, SimSwap reconstructs the entire face, overwriting original pixels and disrupting spatial continuity, which leads to high BER for shallow feature based methods. In contrast, our approach enhances watermark redundancy through DT-CWT in directional sub-bands and leverages SC-GNN to model structural relationships between regions, enabling robust watermark recovery even under significant appearance and semantic changes. Although our proposed semi-robust Detector exhibits a slightly lower BER than SepMark under deepfake attacks, the primary objective of the Detector is to distinguish between authentic and forged content. As long as a significant BER gap exists between common distortions and deepfake forgeries, effective detection can be achieved. Our method yields a substantially lower BER under common distortions, thereby enlarging the separation margin between real and fake samples and significantly enhancing discriminability. However, the existing methods have large fluctuations in BER and lack of detection reliability.
TABLE IV: Cross-dataset robustness and visual quality on LFW, FaceForensics++, DFDC, and Celeb-DF v2. Visual quality is reported with PSNR/SSIM, and robustness with BER (%). $\Delta$ denotes the absolute BER difference compared with CelebA-HQ.
128 $\times$ 128
Metric / Distortion LFW FF++ DFDC Celeb-DF v2
PSNR ↑ 46.27 47.65 47.65 47.67
SSIM ↑ 0.9948 0.9957 0.9945 0.9957
Average Common 0.0101% 0.0098% 0.0059% 0.0076%
$\Delta$ $\uparrow$0.0034% $\uparrow$0.0031% $\downarrow$0.0008% $\uparrow$0.0009%
Average Malicious 0.0389% 0.0278% 0.0225% 0.0246%
$\Delta$ $\uparrow$0.0123% $\uparrow$0.0012% $\downarrow$0.0041% $\downarrow$0.0020%
256 $\times$ 256
Metric / Distortion LFW FF++ DFDC Celeb-DF v2
PSNR ↑ 43.11 43.41 43.37 43.39
SSIM ↑ 0.9908 0.9895 0.9897 0.9894
Average Common 0.0036% 0.0017% 0.0028% 0.0038%
$\Delta$ $\uparrow$0.0013% $\downarrow$0.0006% $\uparrow$0.0005% $\uparrow$0.0015%
Average Malicious 0.0035% 0.0058% 0.0045% 0.0061%
$\Delta$ $\downarrow$0.0012% $\uparrow$0.0011% $\downarrow$0.0002% $\uparrow$0.0014%
@@
TABLE V: ablation results for psnr, ssim, and tracer’s bit error rate (%), where ‘×’ denotes non-usage and $\sqrt{}$’ denotes usage.
@@
PSNR $\uparrow$ 45.1478 43.6807 48.0785 42.9015 47.5121
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9417 0.9957
JpegTest 0.0000% 0.0000% 0.0298% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0157% 0.0000% 0.0000%
GaussianBlur 0.0212% 0.0027% 0.0229% 0.0512% 0.0021%
MedianBlur 0.0023% 0.0015% 0.0071% 0.0084% 0.0011%
Dropout 0.0009% 0.0000% 0.0000% 0.0000% 0.0000%
SaltPepper 0.0043% 0.0043% 0.0408% 0.1251% 0.0187%
GaussianNoise 0.0041% 0.0195% 0.5372% 0.0808% 0.0253%
Average 0.0047% 0.0040% 0.0934% 0.0379% 0.0067%
FSRT 0.0825% 0.0232% 0.1984% 0.0813% 0.0785%
SimSwap 0.2164% 0.2139% 0.5991% 0.1905% 0.1637%
CSCS 0.0098% 0.0056% 0.2025% 0.1221% 0.0091%
HifiFace 0.0074% 0.0088% 0.1109% 0.0113% 0.0081%
Uniface 0.0087% 0.0047% 0.1071% 0.0623% 0.0052%
GANimation 0.0015% 0.0000% 0.0138% 0.0013% 0.0000%
StarGAN (Male) 0.0204% 0.0075% 0.1245% 0.0169% 0.0065%
StarGAN (Young) 0.0112% 0.0081% 0.1278% 0.0114% 0.0071%
StarGAN (BlackHair) 0.0103% 0.0094% 0.1138% 0.0121% 0.0032%
StarGAN (BlondHair) 0.0109% 0.0089% 0.1345% 0.0153% 0.0045%
StarGAN (BrownHair) 0.0176% 0.0078% 0.1374% 0.0187% 0.0064%
Average 0.0389% 0.0289% 0.1759% 0.0532% 0.0266%
TABLE VI: ablation results for psnr, ssim, and detector’s bit error rate (%), where ‘×’ denotes non-usage and $\sqrt{}$’ denotes usage.
@@
PSNR $\uparrow$ 45.1478 43.6807 48.0758 42.9015 47.5121
SSIM $\uparrow$ 0.9956 0.9823 0.9955 0.9417 0.9957
JpegTest 0.0000% 0.0000% 0.0174% 0.0000% 0.0000%
Resize 0.0000% 0.0000% 0.0324% 0.0000% 0.0000%
GaussianBlur 0.0028% 0.0037% 0.0064% 0.0074% 0.0032%
MedianBlur 0.0051% 0.0044% 0.0061% 0.0027% 0.0019%